OpenStudy (anonymous):

One 10.0 Ω resistor is wired in series with two 10.0 Ω resistors in parallel. A 45.0 V battery supplies power to the circuit. What is the total current through the circuit? A. 3.0 A B. 1.5 A C. 0.75 A D. 0.33 A

OpenStudy (michele_laino):



OpenStudy (anonymous):

equilvalent resistance is 30?

OpenStudy (michele_laino):

no, it is: \[\Large {R_{EQ}} = 10 + \frac{1}{{\frac{1}{{10}} + \frac{1}{{10}}}} = ...\Omega \]

OpenStudy (anonymous):

oh oops ! so it is 5!

OpenStudy (michele_laino):

are you sure?

OpenStudy (anonymous):

oh oops i did not see the 10 in front! so it is 15!

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

now, the requested current is: I=45/15=...amperes

OpenStudy (anonymous):

3! so our solution is choice A? :O

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thank you!!:D

OpenStudy (michele_laino):

:)